A person pulls a toboggan for a distance of 40.0 m along the snow with a rope directed 20.0° above the snow. The tension in the rope is 67.0 N.

(a) How much work is done on the toboggan by the tension force?
_________J

(b) How much work is done if the same tension is directed parallel to the snow?
_________J
Answer #1

workdone by tension

Wt = T d cos x

Wt = 67* 40 * cos 20

Wt = 2518.376 J

=======

Wt' = 67 * 40 = 2680 J

======

Comment in case any doubt, will reply for sure.. Goodluck